OpenStudy (anonymous):

On July 2, Natalie deposited $4500 into a savings account that pays 5.5% interest compounded daily. On July 24, how much interest has she earned on principal? Any help at all would be awesome! I have no idea how to do this! Thank you!

OpenStudy (johnweldon1993):

So lets see, interest would look like \(\large V = P(1 + \frac{r}{n})^{nt}\) We know P = 4500, r(in decimal form) = 0.055, n = 365 (it's daily) and t(in years) is (22/365) *it was only in there for 22 days..so the fraction of a year would be 22/365 for 't' So lets plug everything in \[\large V = 4500(1 + \frac{0.055}{365})^{365\times \frac{22}{365}}\] \[\large V = 4500(1 + \frac{0.055}{365})^{22}\] What do you get for V?

OpenStudy (anonymous):

I got 4,514.94

OpenStudy (johnweldon1993):

What I arrived at as well..meaning the total interest earned in those 22 days was $14.94

OpenStudy (anonymous):

How did you get $14.94?

OpenStudy (johnweldon1993):

They want the total interest earned What we just solved for was the total value in the bank at the end of the 22 days So in interest it would be that value MINUS the initial value 4500

OpenStudy (anonymous):

OH! Oh my gosh. That makes total sense. Thank you so much

OpenStudy (johnweldon1993):

Not a problem :)